# Resource Management and Circuit Scheduling for Distributed Quantum Computing Interconnect Networks

### v3

@@ III-A Dynamic Batch-QCirc Scheduling
The workflow for the proposed batch scheduling method is illustrated in Fig. 4, and the corresponding algorithm pseudocode is outlined in Algorithm 1. In each scheduling cycle, a batch of circuits is selected from $\mathcal{QC}$. This batch is denoted by =$\mathcal{B}=\{\text{QCirc}_{m}\mid m\in\mathcal{I}\}$, where $\mathcal{I}\subseteq\{1,\dots,M\}$ is the set of indices corresponding to the selected circuits. The batch size, $S=|\mathcal{I}|$, may vary across scheduling cycles.
@@
We adopt a simple batch selection process, where elements of $\mathcal{QC}$ are sequentially added to $\mathcal{B}$ until the total required number of qubits, $c_{\rm req}=\sum_{m\in\mathcal{I}}{w_{m}}$, exceeds a threshold $\beta c_{\rm tot}$. Here, $c_{\rm tot}$ denotes the total qubit capacity of all available QPUs, given by $c_{\rm tot}=\sum_{j}n_{j}$. The parameter $\beta$, fixed in the range $10<\beta<$ 1, is chosen to ensure that, with high probability, all circuits in the selected batch can be successfully assigned to available QPUs.
@@
The selected batch then undergoes a batch assignment process, referred to as AssignBatch in Algorithm 1, where the circuits are mapped to the available QPUs. We have developed a heuristic batch assignment algorithm that leverages a MILP formulation, which will be described in detail in Sec. IV.
After the batch assignment, any circuits in the batch that remain unassigned are handled immediately using the HandleOverflow procedure, described in Algorithm 2. The proposed overflow handling strategy is inspired by the scheduling method introduced in [8]. As shown in Algorithm 2, the circuits are first sorted according to their qubit requirements $w_{m}$, while the QPUs are sorted based on their expected busy time (EBT). The circuits are then assigned sequentially to the sorted QPUs, ensuring that they are served as soon as resources become available. Notably, since $\beta$ is selected to make overflow events rare, the overflow queue serves primarily as a failsafe rather than a central component of the workflow. Although our framework includes this mechanism for completeness, its use is expected to be minimal in practice.
@@
To improve QPU utilisation, we incorporate a filling mechanism to address idle, unallocated QPUs. The FillQPU algorithm, described in Algorithm 3, selects circuits from $\mathcal{QC}$ and assigns each of them to idle QPUs individually, using a MILP-based optimisation strategy. To minimise the risk of excessive remote gate generation, only circuits with a $\nu_{m}$ value below a specified threshold are considered. The MILP-based Single QCirc assignment process (QCircAssign in Algorithm 3) will be will be described in detail in Sec. V.
@@
Once all QPUs are allocated, the algorithm waits until a fraction $\alpha$ of the total qubit capacity across all QPUs, i.e., $\sum_{j}{N_{j}}$, becomes available, then proceeds to the next scheduling cycle. This mechanism is referred to as TriggerNextCycle in Algorithm 1.
@@
1:Input: $\mathcal{QP}$, $\mathcal{QC}$
@@
4:     batchMap $\leftarrow$ AssignBatch(batch, $\mathcal{QP}$)
@@
1:Input: batch, batchMap, $\mathcal{QP}$
2:Sort unassigned QCircs in batch by decreasing $w$
3:Sort QPUs in $\mathcal{QP}$ by increasing EBT
4:for $\text{QCirc}_{m}$ in sorted QCircs  do
5:     reqQubits $\leftarrow w_{m}$
6:     while reqQubits ¿ 0 do
7:         Assign QCirc to next QPU from sorted $\mathcal{QP}$
8:         Update reqQubits
9:     end while
10:end for
@@
1:Input: $\mathcal{QC}$, $\mathcal{QP}$
@@
5:         if $\nu_{m}\leq\zeta$ and $w_{m}\leq$ avlCapacity then
@@
Algorithm 4 Single-QCirc Scheduler
1:Input: $\mathcal{QC}$, $\mathcal{QP}$
2:while not IsEmpty($\mathcal{QC}$do
3:     for each $\text{QCirc}_{m}$ in $\mathcal{QC}$ do
4:         AssignQCirc($\text{QCirc}_{m}$)
5:         RemoveAssigned($\text{QCirc}_{m}$)
6:     end for
7:end while
III-B Dynamic Single-QCirc Scheduling
Another approach for scheduling QCircs is to schedule them individually. To optimise this assignment, a MILP-based Single-QCirc assignment approach (Sec. V), similar to the one used in the Fill Algorithm (Algorithm 3), is employed.
IV Batch-QCirc Assignment
Batch-QCirc assignment refers to the mapping of a batch of QCircs to available QPUs within the dynamic batch-QCirc scheduling framework. We propose an optimisation algorithm that improves this assignment by targeting three key objectives:
(a) Minimising decoherence-induced infidelity during remote gate executions across all QCircs: Remote gate operations typically incur significantly higher latency than local gates due to the need for both quantum and classical communication across QPUs. This increased latency exposes the computing qubits assigned to the QCircs to greater decoherence. When remote gates are invoked repeatedly within a QCirc, the cumulative effect can result in substantial fidelity loss, thereby degrading the accuracy of distributed quantum computations. Minimising this detrimental impact, considering the effects across the entire batch of QCircs, is therefore one of our primary objectives.
(b) Minimising infidelity from remote entanglement generation: Remote entanglement generation is typically imperfect, introducing some level of infidelity during its implementation. This fidelity can be further altered through operations such as entanglement swapping and entanglement distillation. As discussed in Section 2, we adopt an abstract network model in this paper, where each inter-QPU link is characterised by a fixed infidelity parameter, without delving into the specifics of the underlying physical infrastructure. Similar to the decoherence effects discussed earlier, repeated use of remote gates within a QCirc leads to cumulative fidelity degradation. Our objective is to mitigate this degradation, taking into account the fidelity impact across the entire batch of QCircs.
(c) Maximising the number of assigned QCircs: This objective is primarily aimed at enhancing system throughput and optimising resource utilisation.
The output of the Batch-QCirc assignment process comprises three key components: (a) the partition number assigned to each QCirc, (b) the specific QPUs allocated to each QCirc, and (c) the number of QPU qubits allocated to each circuit partition. This information is subsequently used by the circuit partitioning algorithm to efficiently decompose each QCirc into smaller sub-circuits.
IV-A Problem Formulation
To mathematically formulate the problem of batch-QCirc assignment, we begin by defining the following variables, which represent the outputs we aim to determine:
  • Let $x_{mj}$ be a binary variable representing the number of qubits from ${\rm{QCirc}}_{m}$ that are assigned to ${\rm{QPU}}_{j}$.
  • Let $r_{mj}$ be a binary variable, where $r_{mj}=1$ if $x_{mj}>0$, and $r_{mj}=0$ otherwise.
  • Let $b_{m}$ be a binary variable, where $b_{m}=1$ if ${\rm{QCirc}}_{m}$ is successfully assigned, i.e., $\sum_{j}{r_{mj}}>0$, and $b_{m}=0$ otherwise.
In the following subsections, we formulate the objective functions and constraints to establish our multi-objective optimisation problem.
IV-A1 Objective 1: minimising decoherence-induced infidelity
@@
We define a cost function that captures the degrading impact of qubit decoherence during remote gate executions, as follows:
$C_{\rm dec}=(\frac{1}{T_{\rm dec}})\sum_{m}{\sum_{j_{1}}{\sum_{j_{2}>j_{1}}{\nu_{m}w_{m}r_{mj_{1}}r_{mj_{2}}T_{j_{1}j_{2}}}}},$ (1)
where $T_{j_{1}j_{2}}$ represents the latency parameter associated with a remote gate involving ${\rm{QPU}}_{j_{1}}$ and ${\rm{QPU}}_{j_{2}}$. The term $T_{j_{1}j_{2}}/T_{\rm dec}$ captures the relative decoherence impact of a single remote gate execution. To account for circuit-specific features, we incorporate several weighting factors into the cost function. The qubit count $w_{m}$ reflects the potential exposure of circuit $\text{QCirc}_{m}$ to decoherence. The parameter $\nu_{m}$, defined as
$\nu_{m}=\lambda g_{m}+(1-\lambda)\sigma_{m},\quad 0<\lambda<1,$
effectively captures the internal connectivity pattern of $\text{QCirc}_{m}$, which is highly correlated with the probability of generating remote gates after partitioning. Since the exact number of remote gates per circuit is unknown until after circuit partitioning (which follows batch assignment), we incorporate $\nu_{m}$ into the cost function to represent the circuit’s potential for remote gate generation. This surrogate cost term enables the heuristic optimisation algorithm to make informed scheduling decisions while maintaining computational tractability.
IV-A2 Objective 2: minimising infidelity from remote entanglement generation
The second objective is defined as follows:
$C_{\rm reg}=\sum_{m}{\sum_{j_{1}}{\sum_{j_{2}>j_{1}}{\nu_{m}r_{mj_{1}}r_{mj_{2}}(1-F_{j_{1}j_{2}})}}}.$ (2)
Here, $F_{j_{1}j_{2}}$ represents the fidelity parameter associated with remote entanglement generation between ${\rm{QPU}}{j_{1}}$ and ${\rm{QPU}}{j_{2}}$. Similar to Objective 1, the parameter $\nu_{m}$ is incorporated to account for the circuit’s connectivity pattern and its tendency to produce remote gates.
IV-A3 Objective 3: maximising the number of assigned QCircs
The objective of maximising successfully assigned QCircs can be formulated as:
$C_{\text{asg}}=-\sum_{m=1}^{M}{b_{m}}$ (3)
The negative sign converts the maximisation of assigned QCircs into an equivalent minimisation goal.
IV-A4 Multi-objective Optimisation problem
Our multi-objective optimisation is shown in Problem Formulation 1. Four primary constraints are considered and mathematically represented within this formulation. Their corresponding explanations are provided below.
  • Each quantum circuit, ${\rm{QCirc}}_{m}$, must be either fully assigned or left entirely unassigned.
  • Circuit partitions cannot exceed the qubit capacity of their allocated QPUs.
  • The number of partitions for the quantum circuit ${\rm{QCirc}}_{m}$ is restricted to be less than or equal to $K_{m}^{\max}$, a predefined threshold. This threshold can be specified by the user or the network controller based on various factors, such as quantum data centre policies, circuit size, and error tolerance.
  • Only a single circuit partition is assigend to a QPU.
IV-B MILP-based optimisation method
To formulate our optimisation problem as a MILP, we first linearise the non-linear terms by introducing an auxiliary binary variable [26]:
$z_{mj_{1}j_{2}}=r_{mj_{1}}r_{mj_{2}}.$ (4)
and the following constraints:
$r_{mj_{1}}+r_{mj_{2}}-1\leq z_{mj_{1}j_{2}}\leq\min(r_{mj_{1}},r_{mj_{2}})$ (5)
With this linearisation, the problem can be reformulated as a MILP, as outlined in Problem Formulation 2. The parameter $\gamma$ is a fixed constant larger than $\max\{N_{j}\}$.
To address the multi-objective nature of the problem, we employ the $\varepsilon$-constraint method [27], which converts the problem into a single-objective formulation. Specifically, we minimise $C_{\rm dec}$ while imposing constraints on $C_{\rm reg}$ and $C_{\rm asg}$.
V Single-QCirc Assignment
Single-QCirc assignment focuses on optimally mapping a single QCirc to the available QPUs. To this end, we adapt the MILP formulation introduced in the previous section, simplifying it for the single-circuit case. The resulting MILP formulation is presented in Problem Formulation 3.
@@ V Single-QCirc Assignment
$n_{j}$ Equals to $N_{j}$ if ${\rm{QPU}}_{j}$ is available, and 0 otherwise
@@
$\nu_{m}$ Average number of two-qubit gates acting on each qubit in ${\rm{QCirc}}_{m}$
$x_{mj}$ Number of computing qubits within ${\rm{QPU}}_{j}$ assigned to ${\rm{QCirc}}_{m}$
$r_{mj}$ Equals to 1 if $x_{mj}>0$, and 0 otherwise
$b_{m}$ Equals to 1 if ${\rm{QCirc}}_{m}$ is assigned, and 0 otherwise
@@
$T_{j_{1}j_{2}}$ The latency parameter associated with remote gate execution, involving ${\rm{QPU}}_{j_{1}}$ and ${\rm{QPU}}_{j_{2}}$
@@
$F_{j_{1}j_{2}}$ The fidelity parameter associated with remote remote entanglement generation between ${\rm{QPU}}_{j_{1}}$ and ${\rm{QPU}}_{j_{2}}$
{mdframed}
[linewidth=0.5pt, roundcorner=5pt]
$\min\{C_{\rm dec},C_{\rm reg},C_{\rm asg}\}$
s.t.
$\sum_{j=1}^{J}{x_{mj}}=b_{m}w_{m}\quad\text{for}\quad m=1,\dots,M$
$\sum_{m=1}^{M}{x_{mj}}\leq n_{j}\quad\text{for}\quad j=1,\dots,J$
$\sum_{j=1}^{J}{r_{mj}}\leq K_{m}^{\rm max}\quad\text{for}\quad m=1,\dots,M$
$\sum_{m=1}^{M}{r_{mj}}\leq 1\quad\text{for}\quad j=1,\dots,J$
Figure 1: Batch-QCirc assignment problem formulation.
{mdframed}
[linewidth=0.5pt, roundcorner=5pt]
$\min~{}(\frac{1}{T_{\rm dec}})\sum_{m}{\sum_{j_{1}}{\sum_{j_{2}>j_{1}}{\nu_{m}w_{m}z_{mj_{1}j_{2}}T_{j_{1}j_{2}}}}}$
s.t.
$\sum_{m}{\sum_{j_{1}}{\sum_{j_{2}>j_{1}}{\nu_{m}z_{mj_{1}j_{2}}(1-F_{j_{1}j_{2}})}}}\leq th$
$-\sum_{m=1}^{M}{b_{m}}\leq\kappa$
$\sum_{j=1}^{J}{x_{mj}}=b_{m}w_{m}\quad\text{for}\quad m=1,{\scriptstyle\ldots},M$
$\sum_{m=1}^{M}{x_{mj}}\leq n_{j}\quad\text{for}\quad j=1,{\scriptstyle\ldots},J$
$\sum_{j=1}^{J}{r_{mj}}\leq K_{m}^{\rm max}\quad\text{for}\quad m=1,{\scriptstyle\ldots},M$
$\sum_{m=1}^{M}{r_{mj}}\leq 1\quad\text{for}\quad j=1,{\scriptstyle\ldots},J$
$r_{mj}\leq x_{mj}\leq\gamma r_{mj}\;\;\text{for}\;\;m=1,{\scriptstyle\ldots},M\;\;j=1,{\scriptstyle\ldots},J$
$b_{m}\leq\sum_{j=1}^{J}{r_{mj}}\leq\gamma b_{m}\quad\text{for}\quad m=1,{\scriptstyle\ldots},M$
$z_{mj_{1}j_{2}}\leq r_{mj_{1}}\;\text{for}\;m=1,{\scriptstyle\ldots},M\;j_{1},j_{2}=1,{\scriptstyle\ldots},J\;j_{2}>j_{1}$
$z_{mj_{1}j_{2}}\leq r_{mj_{2}}\;\text{for}\;m=1,{\scriptstyle\ldots},M\;j_{1},j_{2}=1,{\scriptstyle\ldots},J\;j_{2}>j_{1}$
$z_{mj_{1}j_{2}}\geq r_{mj_{1}}+r_{mj_{2}}-1\quad\text{for}\quad m=1,{\scriptstyle\ldots},M$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad j_{1},j_{2}=1,{\scriptstyle\ldots},J\;j_{2}>j_{1}$

### v4

@@ III-A Dynamic Batch QCirc Scheduling
The workflow for the proposed batch scheduling method is illustrated in Fig. 5, and the corresponding algorithm pseudocode is outlined in Algorithm 1. In each scheduling cycle, a batch of circuits is selected from $\mathcal{QC}$. This batch is denoted by $\mathcal{B}=\{\text{QCirc}_{m}\mid m\in\mathcal{I}\}$, where $\mathcal{I}\subseteq\{1,\dots,M\}$ is the set of indices corresponding to the selected circuits. The batch size, $S=|\mathcal{I}|$, may vary across scheduling cycles.
@@
We adopt a simple batch selection process, where elements of $\mathcal{QC}$ are sequentially added to $\mathcal{B}$ based on their requests’ arrival time until the total required number of qubits, $c_{\rm req}=\sum_{m\in\mathcal{I}}{w_{m}}$, exceeds a threshold $\beta c_{\rm tot}$. Here, $c_{\rm tot}$ denotes the total qubit capacity of all available QPUs, given by $c_{\rm tot}=\sum_{j}{s_{j}N_{j}}$. The parameter $\beta$, fixed in the range $0<\beta<1$, is chosen to ensure that, with high probability, all circuits in the selected batch can be successfully assigned to available QPUs.
@@
The selected batch then undergoes a batch assignment process, referred to as AssignBatch in Algorithm 1, where the circuits are mapped to the available QPUs. We have developed an MILP-based batch assignment algorithm, which will be described in detail in Sec. III-A1.
After the batch assignment, there might be a few circuits unassigned if the $\beta$ is large, e.g., close to 1, or some circuits are complex to be distributed among available QPUs given the constraints. Any circuits in the batch that remain unassigned are handled after batch assignment using the HandleOverflow procedure, described in Algorithm 2. The QPUs are first sorted in ascending order based on their expected busy time (EBT). The unassigned circuits are then assigned sequentially to the sorted QPUs, while taking into account circuit constraints such as capacity and the maximum number of partitions. The EBT for a given QPU can be estimated from the execution time of circuits already assigned to it or, when such information is not available, from other factors such as circuit connectivity, the number of partitions assigned to the circuit, and historical assignment data. This overflow step acts primarily as a failsafe, as $\beta$ is chosen to make such cases rare.
@@
To improve QPU utilisation, we incorporate a filling mechanism that assigns circuits to idle, unallocated QPUs. The FillQPU algorithm, described in Algorithm 3, selects circuits from $\mathcal{QC}$ and assigns each of them to idle QPUs individually, using a MILP-based optimisation strategy. To reduce the risk of excessive remote gate generation, only circuits with a $\nu_{m}$ value below a specified threshold are considered. Here, $\nu_{m}$ (also see in Sec. III-A1) is a score parameter that quantifies the circuit’s connectivity and its likelihood of generating costly remote gates in a distributed setting. The MILP-based Single QCirc assignment process (AssignQcirc in Algorithm 3) is detailed in Sec. III-B.
@@
Once all QPUs are allocated, the algorithm waits until a fraction $\alpha$ of the total qubit capacity across all QPUs, i.e., $\alpha\sum_{j}{N_{j}}$, becomes available, then proceeds to the next scheduling cycle. This mechanism is referred to as TriggerNextCycle in Algorithm 1.
@@
1:Input: $\mathcal{QP}$, $\mathcal{QC}$, $G$
@@
4:  batchMap $\leftarrow$ AssignBatch(batch, $\mathcal{QP}$,$G$)
@@
1:Input: batch, batchMap, $\mathcal{QP}$, $G$
[…]
4:for $\text{QCirc}_{m}$ in $U$ do
5:  Scan $\mathcal{QP}_{\rm s}$ to find a set $\{\text{QPU}_{j_{k}}\}_{k=1}^{K}$
6: such that $\sum_{k=1}^{K}N_{j_{k}}\geq w_{m}$ and $K\leq K^{\max}_{m}$
7:  Assign $\text{QCirc}_{m}$ to $\{\text{QPU}_{j_{k}}\}_{k=1}^{K}$
8:  $\mathcal{QP}_{\rm s}\leftarrow\mathcal{QP}_{\rm s}\setminus\{\text{QPU}_{j_{k}}\}_{k=1}^{K}\triangleright$ remove allocated QPUs
9:end for
@@
1:Input: $\mathcal{QC}$, $\mathcal{QP}$, $G$
@@
5:   if $\nu_{m}\leq\Gamma$ and $w_{m}\leq$ avlCapacity then
@@
III-A1 Batch-QCirc Assignment
We propose an optimisation algorithm to minimise the detrimental effects of inter-QPU communication, focusing on decoherence-induced errors and fidelity loss from remote entanglement generation. Remote gates incur significantly higher execution time than local gates due to quantum and classical communication delays and loss, increasing qubit exposure to decoherence. Moreover, remote entanglement generation is inherently imperfect, with fidelity that can degrade further through operations such as entanglement swapping and distillation. Repeated use of remote gates amplifies these effects, causing substantial fidelity loss and degrading the overall computational accuracy of distributed quantum computations. Our approach mitigates this degradation for the batch of QCircs as a group. The output of the Batch-QCirc assignment specifies the number of partitions assigned to each QCirc and the QPUs allocated for their execution, which then guide the circuit partitioning process in efficiently decomposing each QCirc into smaller sub-circuits. To mathematically formulate the problem of Batch-QCirc assignment, we begin by defining the following decision variables:
  • Let $r_{mj}$ be a binary variable, where $r_{mj}=1$ if ${\rm QPU}_{j}$ is allocated to ${\rm QCirc}_{m}$, and $r_{mj}=0$ otherwise.
  • Let $y_{mk}$ be a binary variable, where $y_{mk}=1$ if ${\rm QCirc}_{m}$ is partitioned into exactly $k$ parts (i.e., allocated to $k$ distinct QPUs), and $y_{mk}=0$ otherwise.
@@ III-A1 Batch-QCirc Assignment
We define a cost function that captures the impact of qubit decoherence and imperfect entanglement generation during remote gate executions as follows:
$C$ $=\sum_{m\in\mathcal{I}}\sum_{k=1}^{K_{m}^{\rm max}}\sum_{1\leq j_{1}<j_{2}\leq J}\nu_{mk}\;y_{mk}\;r_{mj_{1}}r_{mj_{2}}\big{(}\omega_{0}\frac{w_{m}T_{j_{1}j_{2}}}{T_{\rm dec}}+$
$\qquad\omega_{1}(1-F_{j_{1}j_{2}})\big{)}$ (1)
where $T_{j_{1}j_{2}}$ is the latency parameter, i.e., average time required to establish entanglement, associated with a remote gate between ${\rm QPU}_{j_{1}}$ and ${\rm QPU}_{j_{2}}$, and $F_{j_{1}j_{2}}$ is the fidelity parameter associated with remote entanglement generation for the QPU pair. The term $(w_{m}T_{j_{1}j_{2}})/T_{\rm dec}$ captures the relative decoherence impact, where larger circuits ($w_{m}$) and longer communication delays ($T_{j_{1}j_{2}}$) increase decoherence relative to the characteristic decoherence parameter $T_{\rm dec}$. The parameters $\omega_{0}$ and $\omega_{1}$ are constant weights that reflect the relative importance of decoherence‑induced errors versus fidelity loss from entanglement generation; unless otherwise stated, we set $\omega_{0}=\omega_{1}$.
@@
$s_{j}$ Equals to 1 if ${\rm{QPU}}_{j}$ is available, and 0 otherwise
@@
$K_{m}^{\rm max}$ partition count threshold for ${\text{QCirc}}_{m}$
$r_{mj}$ Equals to 1 if ${\text{QCirc}}_{m}$ is assigned to ${\text{QPU}}_{j}$, and 0 otherwise
$y_{mk}$ Equals to 1 if ${\text{QCirc}}_{m}$ is partitioned to $k$ parts, and 0 otherwise
@@
$T_{j_{1}j_{2}}$ The latency parameter for establishing entanglement for remote gate execution, involving ${\rm{QPU}}_{j_{1}}$ and ${\rm{QPU}}_{j_{2}}$
@@
$F_{j_{1}j_{2}}$ The fidelity parameter associated with remote entanglement generation between ${\rm{QPU}}_{j_{1}}$ and ${\rm{QPU}}_{j_{2}}$
$\nu_{mk}$ Partitioning cost coefficient for ${\text{QCirc}}_{m}$ and a given partition count $k$
$g_{m}$ Average degree of nodes in the graph model of ${\text{QCirc}}_{m}$
$\sigma_{m}$ Std. dev. of node degrees in the graph model of ${\text{QCirc}}_{m}$
$\lambda_{2,m}$ Algebraic connectivity of the graph model of ${\text{QCirc}}_{m}$
$\gamma_{m}$ Weighted density of the graph model of ${\text{QCirc}}_{m}$
To incorporate circuit structure and connectivity patterns into our MILP objective, we introduce a partitioning cost coefficient $\nu_{mk}$, which estimates the remote-gate overhead and thereby enables circuit-structure-aware resource allocation. Each quantum circuit is represented as a weighted undirected graph, where nodes correspond to qubits and edges correspond to two-qubit gates. The weight of each edge is given by the number of occurrences of gates between the respective qubit pair. Based on this representation, $\nu_{mk}$ is defined as the estimated graph cut size under $k$-way partitioning (with partitions of equal or nearly equal size). Although many distinct $k$-partitions are possible, adopting a standard balanced $k$-partition is sufficient to capture a circuit’s connectivity level and its tendency to generate costly remote gates, while avoiding extra complexity.
To efficiently obtain $\nu_{mk}$ for different circuits without performing explicit graph partitioning, we propose an estimation method based on a linear regression model trained on three structural graph features: weighted density, algebraic connectivity, and coefficient of variance. This regression provides a lightweight estimate of cut size across circuit types, thereby making $\nu_{mk}$ tractable to compute in our MILP formulation. The three features are defined as follows:
(i) Weighted density ($\gamma_{m}$) quantifies the average connectivity per qubit pair, computed as
$\gamma_{m}=\frac{\tfrac{1}{2}\sum_{i=1}^{w_{m}}g_{m}^{(i)}}{\binom{w_{m}}{2}}$, (2)
where $g_{m}^{(i)}$ denotes the weighted degree of qubit node $i$. The numerator corresponds to the total edge weight of the circuit graph, i.e., the total number of two-qubit gate occurrences across the circuit.
(ii) Algebraic connectivity ($\lambda_{2,m}$) is the second-smallest eigenvalue of the normalised Laplacian matrix $L_{m}$ of the circuit graph [31]. It measures global connectivity strength; larger values indicate that the graph is harder to partition without cutting many edges.
(iii) Coefficient of variance ($\sigma_{m}/g_{m}$) captures connectivity imbalance across qubits, defined as
$\frac{\sigma_{m}}{g_{m}}=\frac{\sqrt{\frac{1}{w_{m}}\sum_{i=1}^{w_{m}}(g_{m}^{(i)}-g_{m})^{2}}}{\frac{1}{w_{m}}\sum_{i=1}^{w_{m}}g_{m}^{(i)}},$ (3)
where $g_{m}$ and $\sigma_{m}$ are the average and standard deviation of the weighted degrees $\{g_{m}^{(i)}\}$, respecitively.
These features jointly capture complementary aspects of circuit structure: $\gamma_{m}$ reflects the overall connectivity strength, particularly the density of pairwise interactions; $\lambda_{2,m}$ indicates how robustly the graph remains connected under partitioning; and $\sigma_{m}/g_{m}$ captures connectivity imbalance across qubits that may create partitioning bottlenecks. Importantly, these features scale smoothly with circuit size $w_{m}$, enabling the simple but effective linear regression model to learn structural tendencies rather than raw size.
The regression target is the normalised $k$-way cut size:
$\tilde{\nu}_{mk}=\frac{{\rm Cut}_{m}^{(k)}}{\frac{1}{2}\sum_{i=1}^{w_{m}}g_{m}^{(i)}},$ (4)
where ${\rm Cut}_{m}^{(k)}$ is the sum of edge weights for inter-partition edges when circuit $m$ is optimally partitioned into $k$ balanced parts. We then rescale by the total edge weight to recover the estimated partitioning cost coefficient:
$\nu_{mk}\approx\Big{(}{\hat{\chi}}_{0,k}\gamma_{m}+{\hat{\chi}}_{1,k}\lambda_{2,m}+{\hat{\chi}}_{2,k}\tfrac{\sigma_{m}}{g_{m}}+{\hat{\chi}}_{3,k}\Big{)}\cdot\big{(}{\tfrac{1}{2}\sum_{i=1}^{w_{m}}g_{m}^{(i)}}\big{)}.$ (5)
Here, $({\hat{\chi}}_{0,k},{\hat{\chi}}_{1,k},{\hat{\chi}}_{2,k},{\hat{\chi}}_{3,k})$ denote the regression coefficients, which are pre-learned and fitted by minimising the mean squared error (MSE)
$\frac{1}{|\mathcal{D}|}\sum_{d\in\mathcal{D}}\left(\tilde{\nu}_{dk}-\big{(}\hat{\chi}_{0,k}+\hat{\chi}_{1,k}\gamma_{d}+\hat{\chi}_{2,k}\lambda_{2,d}+\hat{\chi}_{3,k}\tfrac{\sigma_{d}}{g_{d}}\big{)}\right)^{2},$ (6)
for a given $k$ partitions. In this expression, $\mathcal{D}$ denotes a set of quantum circuits with varying types and qubit counts, and $\tilde{\nu}_{dk}$ is the normalised $k$-way cut size for circuit $d\in\mathcal{D}$.
Our optimisation problem, which minimises the cost function $C$ in (1), is formulated as the MILP in Formulation III-A1. The cost function is linearised by introducing the auxiliary variables $z_{mj_{1}j_{2}}=r_{mj_{1}}r_{mj_{2}}$ and $x_{mkj_{1}j_{2}}=z_{mj_{1}j_{2}}y_{mk}$. The role of each constraint in Formulation III-A1 is as follows:
(1) Minimum assignment: Enforces a lower bound on the number of circuits scheduled in the current cycle. By default, $\zeta=M$ so that all circuits in the batch are assigned. If the MILP becomes infeasible, $\zeta$ is decremented until a feasible allocation is found.
(2) Capacity requirement: Ensures that the aggregate qubit capacity of the QPUs assigned to ${\rm QCirc}_{m}$ meets or exceeds its qubit requirement $w_{m}$. While $s_{j}$ (QPU availability binary parameter) may vary between scheduling cycles, it is fixed within each cycle and thus treated as a parameter in the MILP.
(3) Single‑partition per QPU: Ensures that each QPU is assigned at most one circuit partition in a scheduling cycle.
(4) Link between binary variables: Provides an MILP‑compatible link between $r_{mj}$ and $y_{mk}$.
(5)–(6) Linearisation of $z_{mj_{1}j_{2}}$: Implements MILP‑compatible linear formulations of the equalities $z_{mj_{1}j_{2}}=r_{mj_{1}}r_{mj_{2}}$ and $x_{mkj_{1}j_{2}}=z_{mj_{1}j_{2}}y_{mk}$.
In the FillQPU algorithm (Algorithm 3), we employ the parameter $\nu_{mk}$ as a scoring metric for circuit connectivity and remote-gate overhead estimation. In particular, we consider the case $k=2$ (with the index $k$ omitted in Algorithm 3 for simplicity), and only circuits with $\nu_{m}$ not exceeding a threshold $\Gamma$ are selected for individual assignment.